Gold has a face centered cubic unit cell and has the density πŸπŸ—. πŸ‘ π’ˆ π’„π’Žβˆ’πŸ‘. Calculate the length along an edge. (𝐴𝑒 = 197 𝑔 π‘šπ‘œπ‘™βˆ’1)



1
Expert's answer
2021-11-16T19:20:32-0500
"\rho=\frac{m}{a^3}\\19.3=\frac{130.85\cdot10^{-23}}{a^3}\\a=4.08\cdot10^{-8}\ cm"